OpenStudy (anonymous):

Which numbers are equivalent to 0.32? Choose all answers that are correct. A. 3.2% B. 32% C.8/25 D.32/10

OpenStudy (anonymous):

@iGreen

OpenStudy (agl202):

0.32 is the same as 32/100 percent means per 100 so 32/100 = 32% 32/100 can be simplified 32= 4 *8 100 = 4* 25 so 32/100 = 4*8/4*25 = 8/25

OpenStudy (agl202):

@Allsia1245 Do u understand how I did it?

OpenStudy (anonymous):

yea

OpenStudy (agl202):

Ok :)

OpenStudy (agl202):

Medal? Thx...